OpenStudy (anonymous):

5-3(w+4)=w-7 please show work.

OpenStudy (anonymous):

5-3w-12 = w-7 5-12+7 = 4w 4w = 14 w = 14/4 w = 7/2

OpenStudy (anonymous):

the answer is supposed to be 0.

OpenStudy (precal):

5-3w-12=w-7 distribute the -3 to (w+4) -7-3w=w-7 combine like terms (5-12), then add 7 to both sides and add 3w to both sides 0=4w divide by 4 both sides 0=w